# Lightweight and Perceptually-Guided Voice Conversion for Electro-Laryngeal Speech

### v1

@@ 1 Introduction
Nearly 200.000 new cases of laryngeal cancer are diagnosed worldwide each year [12]. Many affected patients need to undergo a total laryngectomy, losing their natural voice source. For voice rehabilitation, a significant number of laryngectomees rely on an electro-larynx device [13]. Most EL speech is characterized by constant, monotonic pitch, reduced prosody, and mechanical noise from the buzzer. These characteristics degrade speech naturalness, expressiveness and intelligibility [8]. To improve EL speech, prior work has investigated several approaches, notably VC techniques [16, 21, 32]. VC aims to transform EL speech into a more natural voice by converting its acoustic features to those of HE speech.
@@ 2.2 Model Architecture
The basic model code is adopted from an unofficial implementation of StreamVC [^2] https://github.com/yuval-reshef/StreamVC. For waveform synthesis, the decoder is trained in a GAN setup, where the discriminator is adopted from the HiFi-GAN [19] discriminator. The final model has approximately 30M parameters (18M content-encoder, 4M speaker-encoder, and 8M decoder) and a total size of 123MB.
@@ 3.3 Fine-tuning
In addition to the original loss functions used in StreamVC, we use several perceptually guided losses. The losses are: (i) Perceptual loss (WavLM) using WavLM [6] features as introduced in FINALLY [3]; (ii) Human feedback (HF) loss, also inspired by FINALLY, based on UTMOS [29] (without PESQ); (iii) $F_{0}$ contour loss via log-$F_{0}$ MSE using pitch contours predicted with fast context-base pitch estimator [^3] https://github.com/CNChTu/FCPE(FCPE); (iv) Intelligibility loss (BNF) as introduced in [36] using bottleneck features (BNF) from the Conformer-CTC [10] model. We fine-tuned the model first on the HE portion of the ELHE database. BNFs are taken from the final layer before the softmax activation; and (v) Intelligibility loss (WEO) using WEOs from Whisper-small instead of BNFs. The Whisper model is also fine-tuned on the HE portion of the database.
@@ 4.1 Objective Evaluation
We first compare different model loss configurations objectively. The results are reported against the source EL and, where possible, to the GT. The evaluation is carried out on a set of 160 utterances.
In addition to our own variants, we use two offline state-of-the-art VC models as baselines: FreeVC [20], and XVC [11]. Since these models are self-supervised, we fine-tune them on the HE portion of the ELHE corpus and evaluate them one-shot on EL inputs against our model. We report results across the objective metrics: (i) Character error rate (CER): Transcriptions are done using zero-shot Whisper-small. CER reflects intelligibility. (ii) DNSMOS [27]: Objective quality predictor model for speech quality (SIG), background noise (BAK), and overall perceptual quality (OVRL). (iii) wvMOS [1]: MOS score predicted by wav2vec2.0. (iv) Speaker similarity (SIM): Speaker similarity calculated via Resemblyzer [30] between converted and target HE speech. (vi) Log-$\bm{F_{0}}$ RMSE: Difference of pitch contour between conversions and target HE speech.
Additionally, we assess robustness under noisy conditions. Random quasi-stationary noise (wind, fan, street traffic, train, waterfalls) and non-stationary (background conversations, sirens, smartphone ringing, in-car sirens) are added to the EL input at varying SNR (0-25dB) levels. CER is computed on the VC speech.
@@ 5 Results & Discussion
Objective results (Table 1, Fig. 3(a)). The overall best objective performance is achieved with +WavLM+HF, with +WEO+HF performing similarly well, while +BNF+HF improves wvMOS and SIM but clearly degrades CER. A general trend is that adding more than 2 auxiliary losses to the GAN setup does not help. The best results come from combining a single perceptual loss (WavLM or WEO) with one speech quality loss (HF). Additional terms (e.g., $F_{0}$ or further intelligibility losses) tend to reduce both intelligibility and quality, likely due to competing gradients preventing convergence. The different losses emphasize distinct properties. WavLM and WEO features capture broader acoustic structure (including low-level noise), which slightly decreases DNSMOS and wvMOS but achieves best CER. By contrast, BNF loss, based on final pre-softmax layer representations, is strongly content-driven and sensitive to imperfect EL-HE alignment. This encourages “mumbled” realizations and grammatical artifacts, explaining the drop in intelligibility despite gains in naturalness and similarity.

### v2

@@ 1 Introduction
Nearly 200.000 new cases of laryngeal cancer are diagnosed worldwide each year [12]. Many affected patients need to undergo a total laryngectomy, losing their natural voice source. For voice rehabilitation, a significant number of laryngectomees rely on an electro-larynx device [13]. Most EL speech is characterized by constant, monotonic pitch, reduced prosody, and mechanical noise from the buzzer. These characteristics degrade speech naturalness, expressiveness and intelligibility [9]. To improve EL speech, prior work has investigated several approaches, notably VC techniques [20, 33, 21, 22]. VC aims to transform EL speech into a more natural voice by converting its acoustic features to those of HE speech.
@@ 2.2 Model Architecture
The model code is adopted from an unofficial implementation of StreamVC [^2] https://github.com/yuval-reshef/StreamVC; all other components were implemented by us. Waveform synthesis is trained in a GAN framework with a discriminator architecture following the HiFi-GAN [18] MPD and MSD design. The final model has approximately 30M parameters (18M content-encoder, 4M speaker-encoder, and 8M decoder) and a total size of 123MB.
@@ 3.3 Fine-tuning
In addition to the original loss functions used in StreamVC, we use several perceptually guided losses. The losses are: (i) Perceptual loss (WavLM) computed on WavLM [6] features as introduced in FINALLY [3]; (ii) Human feedback (HF) loss inspired by FINALLY, defined as the negative mean UTMOS score [30] (without PESQ); (iii) $F_{0}$ contour loss via log-$F_{0}$ MSE using pitch contours predicted with fast context-base pitch estimator [^3] https://github.com/CNChTu/FCPE(FCPE); (iv) Intelligibility loss (BNF) introduced in [37] using MSE between pre-softmax Conformer-CTC [10] bottleneck features (model fine-tuned on the HE portion of ELHE; features taken from the final layer before softmax); and (v) Intelligibility loss (WEO) using MSE between Whisper-small encoder hidden states (Whisper fine-tuned on the HE portion of the database).
@@ 4.1 Objective Evaluation
We first compare different model loss configurations objectively. The results are reported against the source EL and, where possible, to the GT. The evaluation is carried out on a set of 160 unmodified utterances (no PSOLA/DTW). In addition to our own variants, we use two offline state-of-the-art VC models as baselines: FreeVC [19], and XVC [11]. Since no standardized open-source EL-HE VC baselines exist, these models serve as reference points for indirect comparison and are fine-tuned on the HE portion of the ELHE corpus before one-shot evaluated on EL inputs. We report results across the objective metrics: (i) Character error rate (CER): Transcriptions are done using zero-shot Whisper-small. CER reflects intelligibility. (ii) DNSMOS [28]: Objective quality predictor model for speech quality (SIG), background noise (BAK), and overall perceptual quality (OVRL). (iii) wvMOS [1]: MOS score predicted by wav2vec2.0. (iv) Speaker similarity (SIM): Speaker similarity calculated via Resemblyzer [31] between converted and target HE speech. (vi) Log-$\bm{F_{0}}$ RMSE: Difference of pitch contour between conversions and target HE speech. Additionally, we assess robustness under noisy conditions. Random quasi-stationary noise (wind, fan, street traffic, train, waterfalls) and non-stationary (background conversations, sirens, smartphone ringing, in-car sirens) are added to the EL input at varying SNR (0-25dB) levels. CER is computed on the VC speech.
@@ 5 Results & Discussion
Objective results (Table 1, Fig. 3(a)). The overall best objective performance is achieved with +WavLM+HF, with +WEO+HF performing similarly well, while +BNF+HF improves wvMOS and SIM but clearly degrades CER. A general trend is that adding more than 2 auxiliary losses to the GAN setup does not help. The best results come from combining a single perceptual loss (WavLM or WEO) with one speech quality loss (HF). Additional terms (e.g., $F_{0}$ or further intelligibility losses) tend to reduce both intelligibility and quality, likely due to competing gradients preventing convergence. The different losses emphasize distinct properties. WavLM and WEO losses encourage noise-robust phonetic representations that suppress EL-specific buzzing artifacts, improving intelligibility (CER) at the cost of slightly lower perceptual quality scores (DNSMOS, wvMOS). In contrast, the BNF loss, is highly content-constrained and sensitive to EL–HE misalignment, which can over-penalize local timing errors and lead to unstable or “mumbled” realizations, explaining the reduced intelligibility despite gains in naturalness and speaker similarity.